### v1

@@ 3 Random independent sets: a balancing act
As far as we are aware, Algorithm 1 was first given in [57, pp. 244–6], in a slightly more specialized context, which we also discuss later in this chapter.
It remains to prove that Algorithm 1 builds the desired fractional $k$-colouring of $G$. For the analysis, we index the iterations: for $i=0,1,\dots$, let $H_{i}$, ${w}_{i}$ and $\tau_{i}$ denote the corresponding $H$, ${w}$ and $\tau$ in the $i$th iteration of the while loop of Algorithm 1, prior to updating the sequence. Writing $V_{i}=V(H_{i})$ for brevity, we note that $V=V_{0}\supseteq V_{1}\supseteq V_{2}\supseteq\cdots$. We also have ${w}_{i+1}(v)=\sum_{k=0}^{i}\mathbb{P}(v\in X_{V_{k}})\tau_{k}$ for any $v\in V_{i}$, and ${w}_{i+1}(G)=\sum_{k=0}^{i}\tau_{k}$.
@@
then there must be some vertex $u\in V_{i}$ for which ${w}_{i}(u)<1/k$ and ${w}_{i+1}(u)=1/k$, and so $u\in V_{i}\setminus V_{i+1}$ and we are done. We may therefore assume that there is some vertex $u\in V_{i}$ for which $\tau_{i}=(\beta+\gamma\deg(u))/k-{w}_{i}(G)$, and so ${w}_{i+1}(G)=(\beta+\gamma\deg(u))/k$.
For any $j\in\{0,\dots,i\}$, we know by our assumption of the condition in Theorem 3.4, part (a), that
@@
This gives $\beta{w}_{i+1}(u)\geq\beta/k$, and so ${w}_{i+1}(u)=1/k$. This means that $u\in V_{i}\setminus V_{i+1}$, as required. ∎
Later in this chapter, we show how the strongest of the local occupancy conditions, in Definition 3.3, together with some mild additional technical conditions, yields an upper bound of roughly $\beta+\gamma\Delta$ on the (list) chromatic number $\chi(G)$ of $G$.
@@ 4 Local occupancy examples
We begin this section by pointing out that the strongest form of local occupancy holds in the general setting discussed in the Introduction.
@@
Proof.
Let $u$ be a vertex of $G$ and let $F\subseteq G[N(u)]$. Then with $\beta=1+1/\lambda$ and $\gamma=1$, we want to show that $1+\lambda Z^{\prime}_{F}(\lambda)\geq Z_{F}(\lambda)$. But this follows from the fact that $Z_{F}(\lambda)$ is a polynomial with non-negative coefficients such that $Z_{F}(0)=1$. Note that equality holds only when $Z_{F}(\lambda)$ is of degree at most 1. ∎
By Theorem 3.5, we can also conclude from Theorem 4.1 that the hard-core model on $G$ has occupancy fraction at least $1/(1+1/\lambda+\Delta)$, where $\Delta$ is the maximum degree of $G$. As we noted in the Introduction, this bound is attained for the complete graph $K_{\Delta+1}$, and so no choice of $(\beta,\gamma)$ in Theorem 4.1 yields a strictly smaller value of $\beta+\gamma\Delta$.
@@
Further examples of local occupancy require the Lambert $W$-function, which is the inverse of the function $x\mapsto xe^{x}$. We exclusively use the real branch of $W$ defined on $[-1,\infty)$ and require the basic property that $W(x)=\log x-\log\log x+o(1)$ as $x\to\infty$ (see, for example, [25]).
@@
In the limit as $d\to\infty$, provided that $\lambda=o(1)$ with $\log(1/\lambda)=o(\log d)$, such as when $\lambda=1/\log d$, we have $\beta+\gamma d\sim d/\log d$.
It is important to note that the expression for $\beta+\gamma d$ as above is not monotone in $\lambda$. If we are interested in the best asymptotic value as $d\to\infty$, then we are forced to consider $\lambda\to 0$ so that the factor $\frac{1+\lambda}{\lambda}\log(1+\lambda)$ is $1-o(1)$, but we also need that $\lambda$ does not vanish so fast so that $W(d\log(1+\lambda))=o(\log d)$. Any choice of $\lambda=o(1)$ with $\log(1/\lambda)=o(\log d)$ suffices, such as $\lambda=1/\log d$.
Proof.
In a triangle-free graph $G$, the neighbourhood of any vertex is an independent set, and so it suffices to verify the condition of Definition 3.3 when $F$ is an edgeless graph on $y$ vertices, $0\leq y\leq\Delta$. In this case, $Z_{F}(\lambda)=(1+\lambda)^{y}$ and $\lambda Z_{F}^{\prime}(\lambda)/Z_{F}(\lambda)=y\lambda/(1+\lambda)$. Thus, the left-hand side of the inequality (9.4) is the smooth function
@@
With $\beta$ as above, given any $d>0$, we now use the same method to find $\gamma$ which minimizes $h(\gamma)=\beta+\gamma d$. We observe that $h$ is convex on the domain $\gamma>0$, and we solve for the stationary point. Write $\theta=\frac{1+\lambda}{\gamma\lambda}\log(1+\lambda)-1$, so that $\theta>-1$ and
@@
The final assertion follows from the fact that $W(x)=\log x-\log\log x+o(1)$ as $x\to\infty$. ∎
We can now deduce the following corollary.
Corollary 4.3
@@
Proof.
Both of these bounds follow from the asymptotic analysis of $\beta+\gamma\Delta$ in Theorem 4.2. We obtain the first bound from Theorem 3.5 and the monotonicity of $\alpha_{G}(\lambda)$ in terms of $\lambda$, and the second bound follows from Theorem 3.6. ∎
@@
The main reason that we are interested in triangle-free graphs is that Theorem 4.1 is tight in disjoint unions of complete graphs, and triangle-freeness is a natural and simple condition that forces a graph to be maximally ‘far’ from a disjoint union of complete graphs, at least locally. In the above analysis, the lack of triangles as subgraphs manifests itself most clearly as the condition that the neighbourhoods are independent sets. There are many ways to weaken this condition that still result in a significant departure from the behaviour of complete graphs, and we highlight a couple of natural ways for which local occupancy yields good results. Both of our settings are efficiently expressed via an upper bound on the average degree of the graphs $F$, which can appear in the condition of Definition 3.3; the triangle-free case corresponds to enforcing the average degree of such graphs $F$ to be 00.
@@
Proof.
Let $X$ be a random independent set in $G$ from the hard-core model at fugacity $\lambda$. For any $u\in V(G)$, the spatial Markov property gives
@@
Recalling (9.3), the second inequality follows from integrating both sides of the first. ∎
@@
As $d\to\infty$, provided that $\lambda=o(\rho)$ with $\log(\rho/\lambda)=o(\log d)$, such as when $\lambda=1/(\rho\log d)$, we have $\beta+\gamma d\sim d/\log(d/\rho)$.
Proof.
It suffices to verify the condition of Definition 3.3 when $F$ is a graph on $y$ vertices, $0\leq y\leq\Delta$, with average degree at most $\rho$. Then, by Theorem 4.4, $Z_{F}(\lambda)\leq(1+\lambda)^{y}$ and $\lambda Z_{F}^{\prime}(\lambda)/Z_{F}(\lambda)\geq y\lambda/(1+\lambda)^{1+\rho}$. Thus, the left-hand side of the inequality (9.4) is at least $g(y)$, a smooth function given by
@@
We proceed as in the triangle-free case, proving that when $\beta,\rho>0$, $g$ is convex on the domain $y\geq 0$. We then solve for the stationary point $y^{*}$, and setting $g(y^{*})=1$ we solve for $\beta$ to obtain the stated function. As before, given any $d>0$, we can find $\gamma$ which minimizes $h(\gamma)=\beta+\gamma d$ with the same method. When $\rho\geq 0$, $h$ is convex on the domain $\gamma>0$ and the given $\gamma$ is the stationary point of $h$. The asymptotic properties of the function $W$ again yield the final statement. ∎
@@
Proof.
If $F$ has $y$ vertices, then its average degree is at most $y-1$. Similarly, $F$ has at most $t$ edges as each one corresponds to a triangle in $G$ containing $u$, and so the average degree of $F$ is at most $2t/y$. Balancing these bounds gives an average degree of at most $(\sqrt{8t+1}-1)/2\leq\sqrt{2t}$. ∎
We can deduce the following corollary.
Corollary 4.7
@@
$\beta+\gamma\Delta\leq(1+o(1))\frac{\Delta}{W((1+o(1))\Delta\lambda)}\leq(1+o(1))\frac{\Delta}{\log(\Delta/\sqrt{1+t})}.\qed$
@@
An independence number bound of the form in Corollary 4.7 was used in [54] to establish a bound on the sphere packing density of the form $\theta(d)\geq\Omega(d2^{-d})$. An analogue of these ideas in the continuous setting is the main advance of [45], resulting in a stronger lower bound on $\theta(d)$ in terms of the leading constant, and deeper problem-specific insights.
@@
Another natural way to generalize the triangle-free condition is to consider $C_{k}$-free graphs (where $C_{k}$ denotes a cycle on $k$ vertices), $k\geq 3$, of maximum degree $\Delta$. It is not especially clear how bounds on the chromatic number in, say, $C_{8}$-free graphs should be related to those in triangle-free graphs, but nonetheless from local occupancy we get similar bounds in the limit as $\Delta\to\infty$.
It is impossible to fully express $C_{k}$-freeness with a condition that applies exclusively in neighbourhoods, but it is certainly true that in $C_{k}$-free graphs the longest path in a neighbourhood can be on at most $k-2$ vertices. This is because if there is a path $P$ on $\ell$ vertices in $G[N(u)]$, then together with $u$ we find a cycle of length $\ell+1$ in $G$ containing $u$. It turns out that this property alone permits good local occupancy parameters.
@@
Let $k\geq 3$. For any $C_{k}$-free graph $G$ and any vertex $u\in V(G)$, the average degree of any subgraph $F\subseteq G[N(u)]$ is at most $k-3$.
Proof.
Since $G$ is $C_{k}$-free the longest path in $F$ has at most $k-3$ edges. A theorem of Erdős and Gallai [34] states that the average degree of $F$ is then at most $k-3$. ∎
We can then deduce the following corollary.
Corollary 4.9
@@
$\beta+\gamma\Delta\leq(1+o(1))\frac{\Delta}{W((1+o(1))\Delta\lambda)}\leq(1+o(1))\frac{\Delta}{\log(\Delta/k)}.\qed$
A more detailed analysis of this method gives a bound for up to $k<\Delta$, but in this range the bounds degrade quickly. As a way of expressing $C_{k}$-freeness, the fidelity of the condition that neighbourhoods are $P_{k-1}$-free (where $P_{k-1}$ denotes a path on $k-1$ vertices) would appear to diminish as $k$ grows. In particular, we do not expect these methods to be particularly sharp in the case that $k=\Omega(\Delta)$. It would be interesting to extend local occupancy in ways that allow us to capture $C_{\Delta}$-freeness, although such methods would need to be somewhat less local.
@@
A longstanding and difficult conjecture of Reed [60] posits that in any graph $G$ of clique number $\omega$ and maximum degree $\Delta$, the chromatic number $\chi$ satisfies $\chi\leq\lceil(\omega+\Delta+1)/2\rceil$. Elegant partial progress on this conjecture was established by Molloy and Reed [57], who proved that this bound holds for $\chi_{f}$ without the rounding up.
@@
As $\lambda\to\infty$, the hard-core model approaches the uniform distribution on maximum independent sets, so this is a local $(\frac{1}{2}(\omega+1),\frac{1}{2})$-occupancy condition of the form in Theorem 3.4(a) in the case $\lambda=\infty$. This setting is interesting because the stronger condition in Definition 3.3 does not hold with these parameters. Indeed, in the unavoidable case that a subgraph $F\subseteq G[N(u)]$ has independence number $1$, we have $\mathbb{P}(u\in X)=0$ and $\mathbb{E}|X\cap N(u)|=1$, which makes the left-hand side of inequality (9.4) equal to $\frac{1}{2}$. This precludes the condition in Definition 3.3, but if the probability that $\alpha(F_{N(u)})=1$ is sufficiently small, then we can still hope for condition (9.6).
@@
We demonstrate another proof of the inequality (9.6) to highlight a difference between the conditions in Definition 3.3 and Theorem 3.4(a) and some flexibility in the local occupancy methods.
@@
This inequality follows from Markov’s inequality, but we point out that it is simply a consequence of the linear constraints that must be satisfied by probabilities: $p_{i}\geq 0$ for all $i$ and $\sum_{i}p_{i}=1$. The bound (9.6) follows provided that $(\omega-1)p_{0}\geq p_{1}$, which we now derive from the spatial Markov property.
The event $A=1$ corresponds to the realizations of $X$ for which $\alpha(F_{N(u)})=1$. Since we are working in the case $\lambda=\infty$, when $A=1$ we must have $|X\cap N(u)|=1$, because $X\cap N(u)$ is distributed according to a maximum independent set in $F_{N(u)}$. For any such realization of $X$, we can remove the unique occupied neighbour of $u$ and then insert $u$ to obtain another maximum independent set $X^{\prime}$ in which $u$ is occupied. So under $X^{\prime}$ instead, we observe that $A=0$. Each such $X^{\prime}$ arises from at most $|F_{N(u)}|$ different realizations of $X$, and when $A=1$ it must be that $|F_{N(u)}|\leq\omega-1$, since the only way that $A=1$ is if $F_{N(u)}$ is a clique in the neighbourhood of $u$. Thus, $(\omega-1)p_{0}\geq p_{1}$.

### v2

@@ 3 Random independent sets: a balancing act
As far as we are aware, Algorithm 1 was first given in [63, pp. 244–246], in a slightly more specialized context, which we also discuss later in this chapter.
It remains to prove that Algorithm 1 builds the desired fractional $k$-colouring of $G$. For the analysis, we index the iterations: for $i=0,1,\dots$, let $H_{i}$, ${w}_{i}$ and $\tau_{i}$ denote the corresponding $H$, w{w} and $\tau$ in the $i$th iteration of the ‘while’ loop of Algorithm 1, prior to updating the sequence. Writing $V_{i}=V(H_{i})$ for brevity, we note that $V=V_{0}\supseteq V_{1}\supseteq V_{2}\supseteq\cdots$. We also have ${w}_{i+1}(v)=\sum_{k=0}^{i}\mathbb{P}(v\in X_{V_{k}})\tau_{k}$ for any $v\in V_{i}$, and ${w}_{i+1}(G)=\sum_{k=0}^{i}\tau_{k}$.
@@
then there must be some vertex $u\in V_{i}$ for which ${w}_{i}(u)<1/k$ and ${w}_{i+1}(u)=1/k$, and so $u\in V_{i}\setminus V_{i+1}$ and the result follows. We may therefore assume that there is some vertex $u\in V_{i}$ for which $\tau_{i}=(\beta+\gamma\deg(u))/k-{w}_{i}(G)$, and so ${w}_{i+1}(G)=(\beta+\gamma\deg(u))/k$.
For any $j\in\{0,1,\dots,i\}$, we know by our assumption of the condition in Theorem 3.2, part (a), that
@@
This gives $\beta{w}_{i+1}(u)\geq\beta/k$, and so ${w}_{i+1}(u)=1/k$. This means that $u\in V_{i}\setminus V_{i+1}$, as required for Theorem 3.4.
Later in this chapter, we show how the strongest of the local occupancy conditions, in the Key Definition, together with some mild additional technical conditions, yields an upper bound of roughly $\beta+\gamma\Delta$ on the (list) chromatic number $\chi(G)$ of $G$.
@@ 4 Local occupancy examples
We begin this section by pointing out that the strongest form of local occupancy holds in the general setting discussed in Section 1.
@@
To see this, let $u$ be a vertex of $G$ and let $F\subseteq G[N(u)]$. Then with $\beta=1+1/\lambda$ and $\gamma=1$, we want to show that $1+\lambda Z^{\prime}_{F}(\lambda)\geq Z_{F}(\lambda)$. But this follows from the fact that $Z_{F}(\lambda)$ is a polynomial with non-negative coefficients such that $Z_{F}(0)=1$, establishing the theorem. Note that equality holds only when $Z_{F}(\lambda)$ is of degree at most 1.
By Theorem 3.3, we can also conclude from Theorem 4.1 that the hard-core model on $G$ has occupancy fraction at least $1/(1+1/\lambda+\Delta)$, where $\Delta$ is the maximum degree of $G$. As we noted in Section 1, this bound is attained for the complete graph $K_{\Delta+1}$, and so no choice of $(\beta,\gamma)$ in Theorem 4.1 yields a strictly smaller value of $\beta+\gamma\Delta$.
@@
Further examples of local occupancy require the Lambert $W$-function, which is the inverse of the function $x\mapsto xe^{x}$. We exclusively use the real branch of $W$ defined on $1,\infty)$ and require the basic property that $W(x)=\log x-\log\log x+o(1)$ as $x\to\infty$ (see, for example, [29]).
@@
In the limit as $d\to\infty$, provided that $\lambda=o(1)$ with $\log(1/\lambda)=o(\log d)$ (such as when $\lambda=1/\log d$), we have $\beta+\gamma d\sim d/\log d$.
It is important to note that the above expression for $\beta+\gamma d$ is not monotone in $\lambda$. If we are interested in the best asymptotic value as $d\to\infty$, then we are forced to consider $\lambda\to 0$ so that the factor $\frac{1+\lambda}{\lambda}\log(1+\lambda)$ is 1o(1)1-o(1), but we also need that $\lambda$ does not vanish so fast so that $W(d\log(1+\lambda))=o(\log d)$. Any choice of $\lambda=o(1)$ with $\log(1/\lambda)=o(\log d)$ suffices, such as $\lambda=1/\log d$.
Let us now see why Theorem 4.2 holds. In a triangle-free graph $G$, the neighbourhood of any vertex is an independent set, and so it suffices to verify the condition of the Key Definition when $F$ is an edgeless graph on $y$ vertices, $0\leq y\leq\Delta$. In this case, $Z_{F}(\lambda)=(1+\lambda)^{y}$ and $\lambda Z_{F}^{\prime}(\lambda)/Z_{F}(\lambda)=y\lambda/(1+\lambda)$. Thus, the left-hand side of the inequality (8.5) is the smooth function
@@
With $\beta$ as above, and given any $d>0$, we now use the same method to find $\gamma$ which minimizes $h(\gamma)=\beta+\gamma d$. We observe that $h$ is convex on the domain $\gamma>0$, and we solve for the stationary point. Write $\theta=\frac{1+\lambda}{\gamma\lambda}\log(1+\lambda)-1$, so that $\theta>-1$ and
@@
The final assertion follows from the fact that $W(x)=\log x-\log\log x+o(1)$ as $x\to\infty$. This concludes the proof of Theorem 4.2.
With this result, we can now deduce the following corollary.
Corollary
@@
Both of these bounds follow from the asymptotic analysis of $\beta+\gamma\Delta$ in Theorem 4.2. We obtain the first bound from Theorem 3.3 and the monotonicity of $\alpha_{G}(\lambda)$ in terms of $\lambda$, and the second bound follows from Theorem 3.4.
@@
The main reason that we are interested in triangle-free graphs is that Theorem 4.1 is tight in disjoint unions of complete graphs, and triangle-freeness is a natural and simple condition that forces a graph to be maximally ‘far’ from a disjoint union of complete graphs, at least locally. In the above analysis, the lack of triangles as subgraphs manifests itself most clearly as the condition that the neighbourhoods are independent sets. There are many ways to weaken this condition that still result in a significant departure from the behaviour of complete graphs, and we highlight a couple of natural ways for which local occupancy yields good results. Both of our settings are efficiently expressed via an upper bound on the average degree of the graphs $F$, which can appear in the condition of the Key Definition; the triangle-free case corresponds to enforcing the average degree of such graphs $F$ to be 0.
@@
To see this result, let $X$ be a random independent set in $G$ from the hard-core model at fugacity $\lambda$. For any $u\in V(G)$, the spatial Markov property gives
@@
Recalling (8.3), the second inequality follows from integrating both sides of the first, establishing Theorem 4.3.
@@
In the limit as $d\to\infty$, provided that $\lambda=o(\rho)$ with $\log(\rho/\lambda)=o(\log d)$ (such as when $\lambda=1/(\rho\log d)$), we have $\beta+\gamma d\sim d/\log(d/\rho)$.
To prove this, it suffices to verify the condition of the Key Definition when $F$ is a graph on $y$ vertices, $0\leq y\leq\Delta$, with average degree at most $\rho$. Then, by Theorem 4.3, $Z_{F}(\lambda)\leq(1+\lambda)^{y}$ and $\lambda Z_{F}^{\prime}(\lambda)/Z_{F}(\lambda)\geq y\lambda/(1+\lambda)^{1+\rho}$. Thus, the left-hand side of the inequality (8.5) is at least $g(y)$, the smooth function given by
@@
We proceed as in the triangle-free case, proving that when $\beta,\rho>0$, $g$ is convex on the domain $y\geq 0$. We then solve for the stationary point $y^{*}$, and on setting $g(y^{*})=1$ we solve for $\beta$ to obtain the stated function. As before, given any $d>0$, we can find $\gamma$ which minimizes $h(\gamma)=\beta+\gamma d$ with the same method. When $\rho\geq 0$, $h$ is convex on the domain $\gamma>0$ and the given $\gamma$ is the stationary point of $h$. The asymptotic properties of the function $W$ again yield the final statement, establishing Theorem 4.4.
@@
To see this, note that, if $F$ has $y$ vertices, then its average degree is at most y1y-1. Similarly, $F$ has at most $t$ edges since each edge corresponds to a triangle in $G$ containing $u$, and so the average degree of $F$ is at most $2t/y$. Balancing these bounds gives an average degree of at most $(\sqrt{8t+1}-1)/2\leq\sqrt{2t}$, as required for Theorem 4.5.
With this result, we can deduce the following corollary from Theorems 3.3 and 3.4.
Corollary
@@
$\beta+\gamma\Delta\leq(1+o(1))\frac{\Delta}{W((1+o(1))\Delta\lambda)}\leq(1+o(1))\frac{\Delta}{\log(\Delta/\sqrt{1+t})}.$
@@
An independence number bound of the form in the above corollary was used in [60] to establish a bound on the sphere packing density of the form $\theta(d)\geq\Omega(d2^{-d})$. An analogue of these ideas in the continuous setting is the main advance of [50], resulting in a stronger lower bound on $\theta(d)$ in terms of the leading constant, and deeper problem-specific insights.
@@
Another natural way to generalize the triangle-free condition is to consider $C_{k}$-free graphs, with $k\geq 3$, of maximum degree $\Delta$ (where $C_{k}$ denotes a cycle on $k$ vertices). It is not especially clear how bounds on the chromatic number in, say, $C_{8}$-free graphs should be related to those in triangle-free graphs, but nonetheless from local occupancy we get similar bounds in the limit as $\Delta\to\infty$.
It is impossible to fully express $C_{k}$-freeness with a condition that applies exclusively in neighbourhoods, but it is certainly true that in $C_{k}$-free graphs the longest path in a neighbourhood can be on at most k2k-2 vertices. This is because if there is a path $P$ on $\ell$ vertices in $G[N(u)]$, then together with $u$ we find a cycle of length $\ell+1$ in $G$ containing $u$. It turns out that this property alone permits good local occupancy parameters.
@@
Let $k\geq 3$. For any $C_{k}$-free graph $G$ and any vertex $u\in V(G)$, the average degree of any subgraph $F\subseteq G[N(u)]$ is at most k3k-3.
To see this, note that the longest path in $F$ has at most k3k-3 edges since $G$ is $C_{k}$-free. A theorem of Erdős and Gallai [39] states that the average degree of $F$ is then at most k3k-3, establishing Theorem 4.6.
With this result, we can deduce the following corollary from Theorems 3.3 and 3.4.
Corollary
@@
$\beta+\gamma\Delta\leq(1+o(1))\frac{\Delta}{W((1+o(1))\Delta\lambda)}\leq(1+o(1))\frac{\Delta}{\log(\Delta/k)}.$
A more detailed analysis of this method gives a bound for up to $k<\Delta$, but in this range the bounds degrade quickly. As a way of expressing $C_{k}$-freeness, the fidelity of the condition that neighbourhoods are $P_{k-1}$-free (where $P_{k-1}$ denotes a path on k1k-1 vertices) would appear to diminish as $k$ grows. In particular, we do not expect these methods to be particularly sharp when $k=\Omega(\Delta)$. It would be interesting to extend local occupancy in ways that allow us to capture $C_{\Delta}$-freeness, although such methods would need to be somewhat less local.
@@
A longstanding and difficult conjecture of Reed [66] posits that for any graph $G$ of clique number $\omega$ and maximum degree $\Delta$, the chromatic number $\chi$ satisfies $\chi\leq\lceil(\omega+\Delta+1)/2\rceil$. Elegant partial progress on this conjecture was established by Molloy and Reed [63], who proved that this bound holds for $\chi_{f}$ without the rounding up.
@@
As $\lambda\to\infty$, the hard-core model approaches the uniform distribution on maximum independent sets, so this is a local $(\frac{1}{2}(\omega+1),\frac{1}{2})$-occupancy condition of the form in Theorem 3.2(a) in the case $\lambda=\infty$. This setting is interesting because the stronger condition in the Key Definition does not hold with these parameters. Indeed, in the unavoidable case that a subgraph $F\subseteq G[N(u)]$ has independence number $1$, we have $\mathbb{P}(u\in X)=0$ and $\mathbb{E}|X\cap N(u)|=1$, which makes the left-hand side of inequality (8.5) equal to $\frac{1}{2}$. This precludes the condition in the Key Definition, but if the probability that $\alpha(F_{N(u)})=1$ is sufficiently small, then we can still hope for condition (8.7).
@@
We demonstrate another proof of the inequality (8.7) to highlight a difference between the conditions in the Key Definition and Theorem 3.2(a) and some flexibility in the local occupancy methods.
@@
This inequality follows from Markov’s inequality, but we point out that it is also simply a consequence of the linear constraints that must be satisfied by probabilities: $p_{i}\geq 0$ for all $i$ and $\sum_{i}p_{i}=1$. The bound (8.7) follows provided that $(\omega-1)p_{0}\geq p_{1}$, which we now derive from the spatial Markov property.
The event $A=1$ corresponds to the realizations of $X$ for which $\alpha(F_{N(u)})=1$. Since we are working in the case $\lambda=\infty$, when $A=1$ we must have $|X\cap N(u)|=1$, because $X\cap N(u)$ is distributed according to a maximum independent set in $F_{N(u)}$. For any such realization of $X$, we can remove the unique occupied neighbour of $u$ and then insert $u$ to obtain another maximum independent set $X^{\prime}$ in which $u$ is occupied. So under $X^{\prime}$ instead, we observe that $A=0$. Each such $X^{\prime}$ arises from at most $|F_{N(u)}|$ different realizations of $X$, and when $A=1$ it must be that $|F_{N(u)}|\leq\omega-1$, since $F_{N(u)}$ must be a clique in the neighbourhood of $u$ if $A=1$. Thus, $(\omega-1)p_{0}\geq p_{1}$.